# Exploring human-SAV interaction using large language models: The impact of psychological ownership and anthropomorphism on user experience

### v1

@@ Abstract
There has been extensive prior work exploring how psychological factors such as anthropomorphism affect the adoption of shared autonomous vehicles (SAVs). However, limited research has been conducted on how prompt strategies in large language model (LLM)-powered SAV User Interfaces (UIs) affect users’ perceptions, experiences, and intentions to adopt such technology. In this work, we investigate how conversational UIs powered by LLMs drive these psychological factors and psychological ownership, the sense of possession a user may come to feel towards an entity or object they may not legally own. We designed four SAV UIs with varying levels of anthropomorphic characteristics and psychological ownership triggers. Quantitative measures of psychological ownership, anthropomorphism, quality of service, disclosure tendency, sentiment of SAV responses, and overall acceptance were collected after participants interacted with each SAV. Qualitative feedback was also gathered regarding the experience of psychological ownership during the interactions. The results indicate that an SAV conversational UI designed to be more anthropomorphic and to induce psychological ownership improved users’ perceptions of the SAV’s human-like qualities and improved the sentiment of responses compared to a control condition. These findings provide practical guidance for designing LLM-based conversational UIs that enhance user experience and adoption of SAVs.
1 Introduction
@@
Over the past few years, extensive research has been conducted to overcome technical, regulatory, and human-centric barriers to the adoption of Shared Autonomous Vehicles (SAVs), bringing them closer to practical reality. Combining shared mobility with automated driving technologies, SAVs are driverless taxi fleets composed entirely of fully autonomous vehicles (Dai et al., 2021). SAVs offer numerous benefits, such as reduced car ownership and traffic congestion, lower environmental impact, decreased travel time costs, and enhanced accessibility for users (Wu et al., 2023; Huda et al., 2023; Pimenta et al., 2024). Practical advancements in SAV technology are already evident: Waymo provides autonomous ride-hailing services to the public in two U.S. cities (Waymo, 2024), while Apollo Go, a Chinese self-driving travel service platform, operates in 11 cities nationwide in China (Robotgo, 2024). Tesla has also announced plans for a Robotaxi, aiming to commence production before 2027 (Tesla, 2024; Hull et al., 2024). These examples demonstrate the practical advancements and transformative potential of SAV technology.
@@
However, technological advancements alone are insufficient without public willingness to adopt SAVs (Zhang et al., 2019). A positive user experience is essential for the widespread adoption of SAV technology (Flohr et al., 2021). The high levels of autonomy associated with SAVs remove direct human control, which can lead to user unease if effective interaction mechanisms are not in place. This lack of control may negatively impact user comfort and overall acceptance of the technology (Flohr et al., 2021; Kang and Kim, 2020; Wang et al., 2025). To bridge this gap, establishing effective interaction mechanisms is critical. Current research tends to focus mainly on survey studies, driving simulations (e.g., driving simulators and virtual reality), and the physical appearance (Zhang et al., 2019; Zou et al., 2021; Xu et al., 2024; Zhang and Patrick Rau, 2023; Wang et al., 2024) of vehicles, while placing limited emphasis on the key aspect of conversational interactions between humans and SAVs, an area that becomes especially critical as autonomy levels increase. Effective communication is a critical component of user comfort and trust in autonomous systems (Ruijten et al., 2018). Conversational interactions, facilitated through natural language processing such as large language models (LLMs), offer a potential solution to alleviate user concerns by providing a sense of communication, understanding, and responsiveness.
@@
Anthropomorphism, the attribution of human-like characteristics to non-human entities (Salles et al., 2020; Wu et al., 2023), is a key factor in human-SAV interaction that could affect user experience and intention to use SAVs. Research indicates that incorporating simplified human-like features, whether in appearance or behavior, can enhance users’ perceived service value, which in turn fosters loyalty and leads to a more positive interaction experience (Wu et al., 2023). However, a greater level of anthropomorphism may decrease human trust and negatively affect human-SAV interaction quality, especially among respondents without vehicle ownership (Wu et al., 2023). Therefore, finding the optimal level of anthropomorphism is essential for improving user acceptance.
@@
As the name indicates, one of the key features of SAV is the sharing aspect, which involves the removal of private ownership. Over the past few decades, consumer behavior has shifted from the mantra “we are what we own” to “we are what we can access” (Merfeld et al., 2019). Psychological ownership (PO), the feeling of possession towards a target, whether legally owned, shared, or abstract, even in the absence of legal ownership (Pierce et al., 2003; Baxter et al., 2015), is becoming increasingly significant for fostering engagement in access-based services. PO plays a vital role in how people interact with digital products and services, enhancing user engagement, satisfaction, and loyalty towards digital platforms (Kučinskas, 2024; Li and Atkinson, 2020; Liang et al., 2024; Moon et al., 2013; Dai et al., 2021). Fostering a sense of PO when interacting with SAVs can make users feel a sense of personal connection to the shared vehicles. It could bridge the gap between private ownership and shared mobility services, facilitating the transition to shared mobility and ultimately increasing user acceptance in SAV systems (Lee et al., 2019; Dai et al., 2021; Orsot-Dessi et al., 2023).
@@
In recent years, LLMs have become viable tools for enhancing user experience in various digital interaction domains. OpenAI’s ChatGPT, one of the most well-known LLMs, has gained significant popularity for its human-like conversational capabilities since its launch in November 2022 (Lei et al., 2023; Du et al., 2023). LLMs capable of facilitating natural interactions are increasingly influencing the field of intelligent vehicles. Du et al. (2023) explored ChatGPT’s potential in intelligent vehicle domains, such as driver assistance, autonomous driving, human-vehicle interactions, shared control, and enhancing the Driving Intelligence Quotient. Implementing LLMs as UIs for communication with SAVs is expected to foster more natural dialogues between users and vehicles, ultimately enhancing user experience, confidence, safety, and comfort (Gao et al., 2023; Du et al., 2023). When integrated into SAVs, LLMs could play a significant role in triggering PO among SAV users and examining the effects of anthropomorphism. Applying strategies of PO and anthropomorphism when designing conversational UIs for SAVs could optimize human-SAV interaction and enhance users’ perceptions and acceptance of SAVs. However, no prior studies have systematically examined how anthropomorphic and PO-driven conversational agents influence user experience and acceptance of SAVs. This paper investigates these effects by tailoring LLM-based UIs to induce PO or anthropomorphism, aiming to identify strategies that enhance user adoption and satisfaction.
@@
This study aims to highlight differences in user experience across four types of SAV UIs designed using different prompting strategies by evaluating factors such as quality of service, disclosure tendency, interaction enjoyment, SAV response sentiment (i.e., polarity and subjectivity), and user acceptance. It demonstrates how strategic prompts can alter user perceptions, providing actionable insights for the development of conversational UIs in SAVs that foster user experience and acceptance. By focusing on the integration of PO and anthropomorphism, this research fills a critical gap in the literature on human-SAV interactions and contributes to a broader understanding of how these elements affect technology adoption.
The remainder of this article is organized as follows: Section 2 presents and describes the relevant studies and hypotheses of this research. Section 3 introduces the methodology applied, including the user study design, SAV UI development, data collection, and data analysis. Section 4 summarizes and discusses the results. Lastly, Section 5 offers concluding remarks, implications, limitations, and suggestions for future work.
2 Literature review
2.1 Psychological ownership
@@
Psychological ownership refers to the feeling that an object or entity is “MINE” (Pierce et al., 2003). This feeling can occur for any objects that are visible, attractive, interesting, and experienced by an individual (Pierce et al., 2003; Baxter et al., 2015). This concept has gained prominence in consumer research as it can be experienced by all types of audiences even in the absence of physical or legal ownership (Thürridl et al., 2020). With the rise of new technologies shifting the focus from tangible products to intangible experiences, psychological ownership has expanded beyond material possessions to encompass mobile technology services (Liang et al., 2024). In the context of car-sharing autonomous vehicles (AVs), where only one customer request (as an individual or a group) is being taken at a time (Narayanan et al., 2020), the customer temporarily takes “ownership” of an AV (i.e., while they are using this SAV service) without legally owning the vehicle. Thus, the concept of psychological ownership could potentially address a known SAV acceptance barrier caused by the absence of physical ownership (Lee et al., 2019). In this study, we define psychological ownership as a sense of ownership where individuals feel that the SAV they are currently using is their own private AV while they using it (Liang et al., 2024).
@@
Psychological ownership has broad implications and leads to many benefits. In marketing research, psychological ownership is claimed to lead to greater happiness (Li and Atkinson, 2020), increased purchase intention and willingness to pay (Kirk and Swain, 2018; Hingston and Whelan, 2024), more favorable attitudes toward domestic brands (Liang et al., 2024), mediate the negative impact on consumer happiness (Zhao et al., 2023) and enhance stewardship behavior towards public goods (Peck et al., 2021). In the context of digital technology, psychological ownership has been shown to improve engagement and satisfaction with digital products (Kučinskas, 2024), foster loyalty to massively multiplayer online role-playing games (Moon et al., 2013), enhance consumers’ valuation of the digital content (Kirk and Swain, 2016), and serve as a sustainable source of competitive advantage by actively involving customers in value co-creation processes (Kirk and Swain, 2018). In the context of SAVs, psychological ownership is found to positively affect behavioral intention to use SAVs (Lee et al., 2019; Dai et al., 2021; Orsot-Dessi et al., 2023).
@@
It is important to understand when and why users experience a sense of psychological ownership – in other words, how psychological ownership can be triggered. Previous researchers claimed that the sense of ownership is rooted in four fundamental human motives (“roots”): (1) effectance motivation (the desire to interact effectively with the environment to achieve desired outcomes, fostering feelings of efficacy and pleasure), (2) self-identity (possessions help define oneself, express identity to others, and maintain self-continuity over time), (3) home (a sense of anchoring in time and space, offering familiarity, comfort, and security), and (4) stimulation (the need for arousal and activation) (Pierce et al., 2001; Pierce and Peck, 2018). These four motives can be seen as the reason for the existence of the psychological state of ownership rather than its cause (Pierce and Peck, 2018).
@@
The development of psychological ownership occurs through three distinct “routes” (i.e., key experiences) – exercise of control, such as physical control over an object (Atasoy and Morewedge, 2018), control over the design process (Baxter et al., 2015; Baxter and Aurisicchio, 2018), or the perceived ability to influence a target’s behavior (Kirk, 2019); intimate knowledge, such as contamination concerns (Hingston and Whelan, 2024), or repeated positive experiences in and with a place or target (Peck et al., 2021); and, investment of the self, such as naming (Stoner et al., 2018), designing (Kirk et al., 2018), customizing (Li and Atkinson, 2020), and voting for (Fuchs et al., 2010) a product (Baxter et al., 2015; Pierce and Peck, 2018; Peck et al., 2021). Beyond these routes, other strategies have been explored to trigger psychological ownership among consumers including fostering positive affect during brand consumption (Thürridl et al., 2020), using emotional narratives that incorporate elements of the three routes (Guo et al., 2025), and employing actionable manipulations, such as placing signs like “welcome to YOUR park” in public spaces (Peck et al., 2021).
Inspired by increasing interest in how different ways of engaging with a target can influence psychological ownership (Peck et al., 2021), this research incorporates elements of the three routes into the designing process of the SAV UI to trigger psychological ownership when users interact with the vehicle and investigate its potential effects on user experience and acceptance of SAVs.
2.2 Anthropomorphism and large language models
@@
Anthropomorphism refers to the attribution of human-like traits, behaviors, or mental states to non-human entities, including objects, animals, brands, and technological artifacts (Salles et al., 2020; Wu et al., 2023). In the context of SAVs, Wu et al. (2023) classified anthropomorphism into two levels: superficial anthropomorphism and deep anthropomorphism. Superficial anthropomorphism represents lower-level dimensions such as physical appearances (e.g., friendly or aggressive), voice tones (e.g., friendly or cool), and vehicle movement patterns (elegant vs. rigid). In contrast, deep anthropomorphism encompasses higher-level dimensions like authenticity (lifelike or artificial), human-likeness (human-like or machine-like), naturalness (natural or synthetic), and awareness (conscious or unconscious).
@@
Superficial anthropomorphism has been shown to bring numerous benefits to human-computer interactions. For instance, it fosters comfort, familiarity, and perceived trustworthiness during interactions, thereby enhancing the overall user experience with SAVs (Wu et al., 2023). Additionally, anthropomorphic features can elicit more positive user responses by increasing a sense of connectedness when interacting with smart objects (Kang and Kim, 2020). It also positively influences users’ attitudes and facilitates the establishment of user–agent relationships (Wu et al., 2023). On the other hand, deep anthropomorphism may reduce the quality of human-SAV interaction, particularly among low-income, low-education male respondents without vehicle ownership (Wu et al., 2023). Conversely, some studies have shown that human-like qualities positively influence users’ perceived friendliness and safety, augmenting trust and overall acceptance of virtual assistant technology (Calahorra-Candao and Martín-de Hoyos, 2024). These contrasting findings suggest that the impact of different levels of anthropomorphism may vary across contexts and user groups. Given these insights, it is important to explore how anthropomorphism can be effectively integrated into SAV interactions and its influence on user experience.
@@
Advancements in artificial intelligence (AI) have led to the development of conversational agents powered by LLMs, such as OpenAI’s ChatGPT. These models can generate human-like text, enabling more natural and engaging interactions (Flohr et al., 2021). In high-level SAVs, which remove direct human control, passengers primarily interact with the vehicle via digital UIs – either conversational UIs or graphical UIs (Ruijten et al., 2018; Flohr et al., 2021). Conversational UIs allow users to communicate with the vehicle using natural language, akin to conversing with a real driver (McTear et al., 2016). However, if these interfaces fail to understand users’ requests or provide incorrect information, it can lead to negative interaction experiences and even serious issues (Følstad and Brandtzaeg, 2020; Sugisaki and Bleiker, 2020; Flohr et al., 2021). One potential solution to enhance the effectiveness of conversational UIs is the adoption of LLMs. Implementing conversational UIs powered by LLMs with anthropomorphic personalities is expected to improve the system’s understanding of user requests and increase user engagement. This integration can lead to an enhanced user experience and greater acceptance of SAVs. Therefore, our study investigates how different anthropomorphic personalities, enabled by LLMs, affect human-SAV interactions.
2.3 User experience factors

### v2

# Exploring human-SAV interaction using LLMs: The impact of psychological factors on user experience

@@ Abstract
There has been extensive prior work exploring how psychological factors such as anthropomorphism affect the adoption of Shared Autonomous Vehicles (SAVs). However, limited research has been conducted on how prompt strategies in large language models (LLM)-powered conversational SAV agents affect users’ perceptions, experiences, and intentions to adopt such technology. In this work, we investigate how conversational SAV agents powered by LLMs drive these psychological factors, such as psychological ownership, the sense of possession a user may come to feel towards an entity or object they may not legally own. We designed four SAV agents with varying levels of anthropomorphic characteristics and psychological ownership triggers. Quantitative measures of psychological ownership, anthropomorphism, quality of service, disclosure tendency, sentiment of SAV responses, and overall acceptance were collected after participants interacted with each SAV. Qualitative feedback was also gathered regarding the experience of psychological ownership during the interactions. The results indicate that an SAV designed to be more anthropomorphic and to induce psychological ownership improved users’ perceptions of the SAV’s human-like qualities, and its responses were perceived as more positive but also more subjective compared to the control conditions. Qualitative findings support established routes to psychological ownership in the SAV context and suggest that the conversational agent’s perceived performance may also influence psychological ownership. Both quantitative and qualitative outcomes highlight the importance of personalization in designing effective SAV interactions. These findings provide practical guidance for designing conversational SAV agents that enhance user experience and adoption.
I Introduction
@@
Over the past few years, extensive research has been conducted to overcome technical, regulatory, and human-centric barriers to the adoption of Shared Autonomous Vehicles (SAVs), bringing them closer to practical reality. Combining shared mobility with automated driving technologies, SAVs are driverless taxi fleets composed entirely of fully autonomous vehicles [1]. SAVs offer numerous benefits, such as reduced car ownership and traffic congestion, lower environmental impact, decreased travel time costs, and enhanced accessibility for users [2, 3, 4]. Practical advancements in SAV technology are already evident: Waymo provides autonomous ride-hailing services to the public in two U.S. cities [5], while Apollo Go, a Chinese self-driving travel service platform, operates in 11 cities nationwide in China [6]. Tesla has also announced plans for a Robotaxi, aiming to commence production before 2027 [7, 8]. These examples demonstrate the practical advancements and transformative potential of SAV technology.
@@
However, technological advancements alone are insufficient without public willingness to adopt SAVs [9]. A positive user experience (UX) is essential for the widespread adoption of SAV technology [10]. The high levels of autonomy associated with SAVs remove direct human control, which can lead to user unease if effective interaction mechanisms are not in place. This lack of control may negatively impact user comfort and overall acceptance of the technology [10, 11, 12]. To bridge this gap, establishing effective interaction mechanisms is important. Current research tends to focus mainly on survey studies, driving simulations (e.g., driving simulators and virtual reality), and the physical appearance [9, 13, 14, 15, 16] of vehicles, while placing limited emphasis on the key aspect of conversational interactions between humans and SAVs, an area that becomes especially critical as autonomy advances. Such interactions, facilitated through natural language processing such as large language models (LLMs), can foster effective communication, user comfort, trust, and understanding in autonomous systems [17]. Notably, LLM-powered conversational interfaces may not only enhance user comfort by enabling natural dialogue but also endow systems with human-like traits, leveraging anthropomorphism to enhance perceived accuracy [18].
@@
Anthropomorphism, the attribution of human-like characteristics to non-human entities [19, 2], is a key factor in human-SAV interaction that could affect user experience and intention to use SAVs. Research indicates that incorporating simplified human-like features, whether in appearance or behavior, can enhance users’ perceived service value, which in turn fosters loyalty and leads to a more positive interaction experience [2]. However, a greater level of anthropomorphism may decrease human trust and negatively affect human-SAV interaction quality, especially among respondents without vehicle ownership [2]. Therefore, finding the optimal level of anthropomorphism is essential for improving user acceptance.
@@
As the name indicates, one of the key features of SAV is the sharing aspect, which involves the removal of private ownership. Over the past few decades, consumer behavior has shifted from the mantra “we are what we own” to “we are what we can access” [20]. Psychological ownership (PO), the feeling of possession towards a target, whether legally owned, shared, or abstract, even in the absence of legal ownership [21, 22], is becoming increasingly significant for fostering engagement in access-based services. PO plays a vital role in how people interact with digital products and services, enhancing user engagement, satisfaction, and loyalty towards digital platforms [23, 24, 25, 26, 1]. Fostering a sense of PO when interacting with SAVs can make users feel a sense of personal connection to the shared vehicles. It could bridge the gap between private ownership and shared mobility services, facilitating the transition to shared mobility and ultimately increasing user acceptance in SAV systems [27, 1, 28].
@@
In recent years, LLMs have become viable tools for enhancing user experience in various digital interaction domains. OpenAI’s ChatGPT, one of the most well-known LLMs, has gained significant popularity for its human-like conversational capabilities since its launch in November 2022 [29, 30]. LLMs capable of facilitating natural interactions are increasingly influencing the field of intelligent vehicles. [30] explored ChatGPT’s potential in intelligent vehicle domains, such as driver assistance, autonomous driving, human-vehicle interactions, shared control, and enhancing the Driving Intelligence Quotient. Implementing LLMs in developing agents for communication with SAVs is expected to foster more natural dialogues between users and vehicles, ultimately enhancing user experience, confidence, safety, and comfort [31, 30]. When integrated into SAVs, LLMs could play a significant role in triggering PO among SAV users and examining the effects of anthropomorphism. Applying strategies of PO and anthropomorphism when designing conversational agents for SAVs could optimize human-SAV interaction and enhance users’ perceptions and acceptance of SAVs. However, no prior studies have systematically examined how anthropomorphic and PO-driven conversational agents influence user experience and acceptance of SAVs. This paper investigates these effects by tailoring LLM-based conversational SAV agents to induce PO or anthropomorphism, aiming to identify strategies that enhance user adoption and satisfaction.
@@
This study aims to highlight differences in user experience across four types of SAV agents designed using different prompting strategies by evaluating factors such as quality of service, disclosure tendency, SAV response sentiment (i.e., polarity and subjectivity), and user acceptance. It demonstrates how strategic prompts can alter user perceptions, providing actionable insights for the development of conversational agents in SAVs that foster user experience and acceptance. By focusing on the integration of PO and anthropomorphism, this research fills a critical gap in the literature on human-SAV interactions and contributes to a broader understanding of how these elements affect technology adoption.
II Literature review
II-A Psychological ownership
@@
Psychological ownership is the feeling that an object or entity is “MINE” [21]. This feeling can occur for any objects that are visible, attractive, interesting, and experienced by an individual [21, 22]. This concept has gained prominence in consumer research as it can be experienced by all types of audiences even in the absence of physical or legal ownership [32]. With the rise of new technologies shifting the focus from tangible products to intangible experiences, psychological ownership has expanded beyond material possessions to encompass mobile technology services [25]. In the context of car-sharing autonomous vehicles (AVs), where only one customer request (as an individual or a group) is being taken at a time [33], the customer temporarily takes “ownership” of an AV (i.e., while they are using this SAV service) without legally owning the vehicle. Thus, the concept of psychological ownership could potentially address a known SAV acceptance barrier caused by the absence of physical ownership [27]. In this study, we define psychological ownership a user feeling that an SAV is “mine” for the duration of the ride [25].
@@
Psychological ownership has broad implications and leads to many benefits. In marketing research, psychological ownership is claimed to lead to greater happiness [24], increased purchase intention and willingness to pay [34, 35], more favorable attitudes toward domestic brands [25], mediate the negative impact on consumer happiness [36] and enhance stewardship behavior towards public goods [37]. In the context of digital technology, psychological ownership has been shown to improve engagement and satisfaction with digital products [23], foster loyalty to massively multiplayer online role-playing games [26], enhance consumers’ valuation of the digital content [38], and serve as a sustainable source of competitive advantage by actively involving customers in value co-creation processes [34]. In the context of SAVs, psychological ownership is found to positively affect behavioral intention to use SAVs [27, 1, 28].
@@
It is important to understand when and why users experience a sense of psychological ownership – in other words, how psychological ownership can be triggered. Previous researchers claimed that the sense of ownership is rooted in four fundamental human motives (“roots”): (1) effectance motivation (the desire to interact effectively with the environment to achieve desired outcomes, fostering feelings of efficacy and pleasure), (2) self-identity (possessions help define oneself, express identity to others, and maintain self-continuity over time), (3) home (a sense of anchoring in time and space, offering familiarity, comfort, and security), and (4) stimulation (the need for arousal and activation) [39, 40]. These four motives can be seen as the reason for the existence of the psychological state of ownership rather than its cause [40].
@@
The development of psychological ownership occurs through three distinct “routes” (i.e., key experiences) – exercise of control, such as physical control over an object [41], control over the design process [22, 42], or the perceived ability to influence a target’s behavior [43]; intimate knowledge, such as contamination concerns [35], or repeated positive experiences in and with a place or target [37]; and, investment of the self, such as naming [44], designing [45], customizing [24], and voting for [46] a product [22, 40, 37]. Beyond these routes, other strategies have been explored to trigger psychological ownership. This includes fostering positive affect during brand consumption [32], using emotional narratives that incorporate elements of the three routes [47], and employing actionable manipulations, such as placing signs like “welcome to YOUR park” in public spaces [37].
II-B Anthropomorphism
@@
Anthropomorphism refers to the attribution of human-like traits, behaviors, or mental states to non-human entities, including objects, animals, brands, and technological artifacts [19, 2]. In the context of SAVs, [2] classified anthropomorphism into two levels: superficial anthropomorphism and deep anthropomorphism. Superficial anthropomorphism represents lower-level dimensions such as physical appearances (e.g., friendly or aggressive), voice tones (e.g., friendly or cool), and vehicle movement patterns (elegant vs. rigid). In contrast, deep anthropomorphism encompasses higher-level dimensions like authenticity (lifelike or artificial), human-likeness (human-like or machine-like), naturalness (natural or synthetic), and awareness (conscious or unconscious).
@@
Superficial anthropomorphism has been shown to bring numerous benefits to human-computer interactions. For instance, it fosters comfort, familiarity, and perceived trustworthiness during interactions, thereby enhancing the overall user experience with SAVs [2]. Additionally, anthropomorphic features can elicit more positive user responses by increasing a sense of connectedness when interacting with smart objects [11]. It also positively influences users’ attitudes and facilitates the establishment of user–agent relationships [2]. On the other hand, deep anthropomorphism may reduce the quality of human-SAV interaction, particularly among low-income, low-education male respondents without vehicle ownership [2]. Conversely, some studies have shown that human-like qualities positively influence users’ perceived friendliness and safety, augmenting trust and overall acceptance of virtual assistant technology [48]. These contrasting findings suggest that the impact of different levels of anthropomorphism may vary across contexts and user groups. Thus, it is important to explore how anthropomorphism can be effectively integrated into SAV interactions and its influence on user experience.
II-C Conversational interfaces and large language models
In high-level SAVs, specifically those at SAE Levels 4 and 5 where direct human control is largely or entirely removed, passengers primarily interact with the vehicles via digital User Interfaces (UIs), which can be conversational UIs, graphical UIs, or more directly via voice comments [17, 10, 49]. Conversational UIs and command-based interfaces allow users to communicate with the vehicle using natural language, akin to conversing with a real driver [50]. However, if these interfaces fail to understand users’ requests or provide incorrect information, it can lead to negative interaction experiences and even serious issues [51, 52, 10]. Another key challenge is developing explainable autonomous vehicle systems to ensure effective human–vehicle communication. Such systems enhance user trust, promote transparency and accountability in the vehicle’s decision-making processes, and contribute to overall safety, user experience, and comfort [49].
@@ II-C Conversational interfaces and large language models
Advancements in artificial intelligence (AI) have led to the development of conversational agents powered by LLMs, such as OpenAI’s ChatGPT. These models can generate human-like text, enabling more natural and engaging interactions [10, 53]. One potential solution to enhance the effectiveness of the conversational ability of SAVs is the adoption of LLMs. Implementing conversational agents powered by LLMs with anthropomorphic personalities is expected to improve the system’s understanding of user requests and increase user engagement. This integration can lead to an enhanced user experience and greater acceptance of SAVs. Additionally, inspired by growing interest in how different forms of engagement influence psychological ownership [37], this study incorporates elements from the three established routes to psychological ownership into the design of conversational SAV agents to trigger users’ sense of ownership during interactions.
Unlike traditional rule-based or menu-driven in-car chatbots, an LLM-powered SAV agent enables open-domain dialogue. This allows passengers to freely ask questions or make requests, such as requesting route changes, asking for local recommendations, or sharing emotions, without being confined to a predefined set of commands. Such flexibility fosters more natural, human-like interactions that adapt to each passenger’s needs and context, better simulating the complexity and emotional nuances of real-world SAV use.
Therefore, this research investigates how different anthropomorphic personalities and psychological ownership strategies, enabled by LLMs, influence human–SAV interactions, user experience, and acceptance.
II-D User experience factors